# Overview of the ESCAPE Dark Matter Test Science Project for Astronomers

### v1

@@ 1 Introduction
After the discovery of surprisingly large velocity dispersions for galaxies in the Coma cluster (Zwicky, 1933), and following the discovery of discrepancies between spiral galaxy rotation curves and their observed stellar masses (Rubin & Ford, 1970), evidence has continued to mount for a missing mass component in the Universe. This dark matter has long been proposed as the explanation for various physical phenomena, and while there have been various possibilities put forward as to its nature (Bertone et al., 2005; Peter, 2012), cold dark matter (CDM) has proven to be the most enduring; such non-relativistic particles have a short free-streaming length than warmer variants, causing them to quickly coalesce under gravity and allowing for the rapid formation of small-scale structures observed in the early Universe (Blumenthal et al., 1984; Frenk & White, 2012; Conselice, 2014).
Dark matter is thought to play a key role in the formation of structure on both galactic and cosmological scales, so to shine a light on its properties requires understanding the dark matter power spectrum and its evolution over cosmic time. Hence, multiple methods are employed to help constrain these properties (Buckley & Peter, 2018; Mayer, 2022). On the largest scales, for example, the power spectrum of Cosmic Microwave Background (CMB) temperature fluctuations is accurately fitted by CDM, with the ratios of peaks providing measures of the relative densities of baryonic and non-baryonic matter in the Universe, among other cosmological parameter constraints. Meanwhile, x-ray emission from galaxy clusters indicates the presence of large quantities of hot gas that can only be explained should the clusters have a large dark matter component (Allen et al., 2002). These dark matter haloes within which galaxies and galaxy clusters reside can also result in cosmic shear that, when modelled, provides constraints on the dark matter density (e.g. To et al., 2021). Additionally, the alignment of galaxies along our line of sight can produce gravitational lenses, where the foreground galaxy acts as a ‘lens’ (albeit without focusing) around which the background galaxy is distorted and magnified; the modelled dark matter haloes of these lenses contain a central ‘cuspy’ slope (e.g. Shajib et al., 2021; Sonnenfeld & Cautun, 2021) and potential substructures that can be used to constrain the properties of dark matter (e.g. Amorisco et al., 2022; He et al., 2022). Despite the success of CDM, there still exists a number of astrophysical problems yet to be fully resolved, primarily arising from simulations of CDM haloes. These include the “core-cusp” problem (Flores & Primack, 1994; Moore, 1994), the “missing satellites” problem (Klypin et al., 1999; Moore et al., 1999) and subsequent “too many satellites” problem (Kelley et al., 2019; Kim & Peter, 2021), and the “too big to fail” problem (Boylan-Kolchin et al., 2011, 2012). As such, refinements to simulations and further observations from new large-scale surveys like the Euclid survey (Laureijs et al., 2011; Euclid Collaboration et al., 2024) and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST, Ivezić et al., 2019) are underway to consolidate the properties of dark matter on astrophysical scales.
@@
Particle physics and astroparticle physics offer another, more local approach to understanding dark matter: numerous experiments in particle physics have been seeking to observe dark matter particles themselves, whether by producing dark matter itself, or by detecting it directly or indirectly through its interactions with ordinary matter. For more comprehensive reviews of dark matter candidates, the astrophysical evidence, and the range of particle physics experiments aiming to find them, see the following: Sumner (2002); Bertone et al. (2005); Peter (2012); Bauer & Plehn (2017); Baudis (2018); Bertone & Tait (2018); Tao (2020); Balazs et al. (2024), along with the road map for the next decade of US research proposed in the latest Snowmass dark matter complementarity report (Boveia et al., 2025). For detailed reviews of the state of particle physics as a whole, see for example, Workman et al. (2022) and Navas et al. (2024). Among the leading candidates for dark matter are weakly interacting massive particles (WIMPs): these theoretical CDM particles only interact through gravity and arise from extensions to the Standard Model (Roszkowski et al., 2018). However, given the above astrophysical problems with CDM and that no clear detections of WIMPs have yet been observed in particle physics experiments, with other theoretical options also being considered (Bertone & Tait, 2018). These theories include alternative CDM candidates such as axion-like particles (Nagano et al., 2019) and primordial black holes (e.g. Jedamzik, 2020; Green & Kavanagh, 2021; Green, 2024), non-CDM particles such as warm dark matter (WDM; e.g., Dayal & Giri, 2024) and self-interacting dark matter (SIDM; Adhikari et al., 2022), as well as modifications to general relativity (e.g. Belgacem et al., 2019).
@@ 1.1 Synergies in dark matter searches
The European Consortium for Astroparticle Theory (EuCAPT; Alves Batista et al., 2021) aims to coordinate ideas, activities, resources and open environments for the European community of theoretical astroparticle physicists and cosmologists. The iDMEu project (the initiative for Dark Matter in Europe and Beyond; Cirelli et al., 2024) aims to provide a common platform to facilitate both cross-community dark matter discussions and the collection of resources in an online meta-repository, supported by ECFA (the European Committee for Future Accelerators), NuPECC (the Nuclear Physics European Collaboration Committee) and APPEC (the Astroparticle Physics European Consortium).

### v2

@@ 1 Introduction
After the discovery of surprisingly large velocity dispersions for galaxies in the Coma cluster (Zwicky, 1933), and following the discovery of discrepancies between spiral galaxy rotation curves and their observed stellar masses (Rubin & Ford, 1970), evidence has continued to mount for a missing mass component in the Universe. This dark matter has long been proposed as the explanation for various physical phenomena, and there have been various possibilities put forward as to its nature (Bertone et al., 2005; Peter, 2012). There are two main types of dark matter: baryonic (primarily non-luminous massive compact halo objects, or MACHOS, such as black holes and brown dwarfs) and non-baryonic (primarily in the form of new particles), with the Universe consisting of a much higher fraction of the latter – the matter density of non-baryonic matter is approximately five times that of all baryonic matter (Planck Collaboration et al., 2020b), and with about 30% of all baryons accounted for through works by e.g. de Graaff et al. (2019) the amount of remaining ’baryonic dark matter‘ implies a ratio of non-baryonic-to-baryonic dark matter of about 7.6. Hence, for the rest of this review we primarily consider only the non-baryonic component. Cold cold dark matter (CDM) has proven to be the most enduring candidate: such non-relativistic particles have a short free-streaming length than warmer variants, causing them to quickly coalesce under gravity and allowing for the rapid formation of small-scale structures observed in the early Universe (Blumenthal et al., 1984; Frenk & White, 2012; Conselice, 2014).
[…]
Despite the success of CDM, there still exists a number of astrophysical problems yet to be fully resolved. For example, another source of astrophysical evidence for dark matter comes from the anomalous absorption of the redshifted 21cm line from the early Universe in data from the Experiment to Detect the Global Epoch of Reionization Signature (EDGES, Bowman et al., 2018). This absorption was found to be over two times greater than their largest predictions, suggestive of significant cooling of primordial hydrogen gas caused by interactions between dark matter and baryons, favouring baryonic dark matter over non-baryonic (Barkana, 2018; McGaugh, 2018). However, this detection has not been subsequently confirmed, such as the non-detection by the Shaped Antenna measurement of the background RAdio Spectrum 3 experiment (SARAS 3, Singh et al., 2022), with other works having sought alternative explanations (see e.g. Bevins et al. (2022) and references therein). Other problems with CDM primarily arise from simulations of CDM haloes. These include the “core-cusp” problem (Flores & Primack, 1994; Moore, 1994), the “missing satellites” problem (Klypin et al., 1999; Moore et al., 1999) and subsequent “too many satellites” problem (Kelley et al., 2019; Kim & Peter, 2021), and the “too big to fail” problem (Boylan-Kolchin et al., 2011, 2012). As such, refinements to simulations and further observations from new large-scale surveys like the Euclid survey (Laureijs et al., 2011; Euclid Collaboration et al., 2024) and the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST, Ivezić et al., 2019) are underway to consolidate the properties of dark matter on astrophysical scales.
@@
Particle physics and astroparticle physics offer another, more local approach to understanding dark matter: numerous experiments in particle physics have been seeking to observe dark matter particles themselves, whether by producing dark matter itself, or by detecting it directly or indirectly through its interactions with ordinary matter. For more comprehensive reviews of dark matter candidates, the astrophysical evidence, and the range of particle physics experiments aiming to find them, see the following: Sumner (2002); Bertone et al. (2005); Peter (2012); Bauer & Plehn (2017); Baudis (2018); Bertone & Tait (2018); Tao (2020); Balazs et al. (2024); Arbey & Mahmoudi (2021); Oks (2021); Cirelli et al. (2024b), along with the road map for the next decade of US research proposed in the latest Snowmass dark matter complementarity report (Boveia et al., 2025). For detailed reviews of the state of particle physics as a whole, see for example, Workman et al. (2022) and Navas et al. (2024). Among the leading candidates for dark matter are weakly interacting massive particles (WIMPs): these theoretical CDM particles only interact through gravity and arise from extensions to the Standard Model (Roszkowski et al., 2018). However, given the above astrophysical problems with CDM and that no clear detections of WIMPs have yet been observed in particle physics experiments, with other theoretical options also being considered (Bertone & Tait, 2018). These theories include alternative CDM candidates such as axion-like particles (Nagano et al., 2019) and primordial black holes (e.g. Jedamzik, 2020; Green & Kavanagh, 2021; Green, 2024), non-CDM particles such as warm dark matter (WDM; e.g., Dayal & Giri, 2024) and self-interacting dark matter (SIDM; Adhikari et al., 2022), as well as modifications to general relativity (e.g. Belgacem et al., 2019).
@@ 1.1 Synergies in dark matter searches
The European Consortium for Astroparticle Theory (EuCAPT; Alves Batista et al., 2021) aims to coordinate ideas, activities, resources and open environments for the European community of theoretical astroparticle physicists and cosmologists. The iDMEu project (the initiative for Dark Matter in Europe and Beyond; Cirelli et al., 2024a) aims to provide a common platform to facilitate both cross-community dark matter discussions and the collection of resources in an online meta-repository, supported by ECFA (the European Committee for Future Accelerators), NuPECC (the Nuclear Physics European Collaboration Committee) and APPEC (the Astroparticle Physics European Consortium).